OpenStudy (anonymous):

Verify the trigonometric identity. cot(x)sec^4(x)=cot(x)+2tan(x)+tan^3(x)

OpenStudy (raden):

sec^4 x = sec^2 x * sec^ 2 x = (sec^2 x)^2

OpenStudy (raden):

cot(x)sec^4(x) = cot(x)(sec^2(x))^2

OpenStudy (raden):

then use the identity : sec^2(x) = tan^2(x) + 1 therefore, the equation above can be cot(x)(sec^2(x))^2 = cot(x)(tan^2(x) + 1)^2 expand the term in paranthesis to get = cot(x) (tan^4(x) + 2 tan^2(x) + 1) simplify again : = cot(x) tan^4(x) + 2cot(x)tan^2(x) + cot(x) =tan^4(x)/tan(x) + 2tan^2(x)/tan(x) + cot(x) = tan^3(x) + 2tan(x) + cot(x) proof

OpenStudy (anonymous):

why when you simplify again so you get the cot and such

OpenStudy (anonymous):

@primeralph

OpenStudy (primeralph):

Hello?

OpenStudy (anonymous):

i got it :) never mind but thankyou @primeralph

OpenStudy (primeralph):

Okay, you're welcome.